OpenStudy (anonymous):

I have already answered it but please check my work. Prove: sin θ - sin θ•cos^2 θ = sin^3 θ.

OpenStudy (anonymous):

cos^2= 1-sin^2 sin-sin(1-sin^2)=sin^3 distribute sin-sin-sin^3=sin^3 sin^3=sin^3

OpenStudy (anonymous):

@jim_thompson5910 @Luigi0210 @Whitemonsterbunny17

OpenStudy (johnweldon1993):

As long as you see the typo above then yes that is correct \(\large cos^2= 1-sin^2\) \(\large sin-sin(1-sin^2)=sin^3\) distribute \(\large sin-sin\color\red{+}sin^3=sin^3\) \(\large sin^3=sin^3\)

OpenStudy (anonymous):

Thank you so much for your help @johnweldon1993